In the coal-gasification process, carbon monoxide reacts with water to produce carbon dioxide and hydrogen gas.

CO (g) + H2O (g) CO2 (g) + H2 (g)

In an experiment, 0.35 mol of CO and 0.40 mol of H2O were placed in a 1.00-L reaction vessel. At equilibrium, there were 0.22 mol of CO remaining. Keq at the temperature of the experiment is ________.

Group of answer choices

5.5

0.28

0.75

3.5

1.0
Answer #1

CO + H2O \rightleftharpoons CO2 + H2

Initial 0.35 0.40 0 0

Equilibrium. 0.35-x 0.40-x x x

Now, 0.35-x = 0.22

x = 0.35-0.22 = 0.13

So, [H2O] = 0.40-0.13 = 0.27

Keq = [CO2][H2]/[CO][H2O]

= 0.13×0.13/0.22×0.27

= 0.28

Answer is 0.28